Joker stock has a sustainable growth rate of 9 percent, ROE of 17 percent, and dividends per share of $1.55. If the P/E ratio is 15.5, what is the value of a share of stock? (Do not round intermediate calculations. Round your answer to 2 decimal places. Omit the "$" sign in your response.)

Share of stock $________
Answer #1

Solution-

Growth (g) = 9%

ROE (r) = 17%

g= b*r

9%= b *17%

b = 9/17

Retention ratio = 9/17

Dividend payout ratio = 8/17

Dividend = $1.55

Earning per share = 1.55*17/8= $3.29375

PE ratio = 15.5

Value of share = 15.5*3.29375= $51.05